A160 ohm resistor is connected in series with a 0.
Answer #1

R = 160 Ω
L = 0.3 H
C = 4.8 uF
e = 20 * sin(377*t)

ω = 377
XL = ω*L = 377 * 0.3 = 113.1 Ω
XC = 1/ω*C = 1/(377 * 4.8 * 10^-6) = 552.6 Ω

Impedance, Z = sqrt(R^2 + (XL - XC)^2)
Z = sqrt(160^2 + (552.6 - 113.1)^2)
Z = 467.7 Ω

Ipeak = Vmax/Z
Ipeak = 20/467.7
Ipeak = 0.0427 A

Irms = Ipeak/sqrt(2)
Irms = 0.0427/sqrt(2)
Irms = 0.030 A